Ques one - We all know that the price of a postage stamp is $ .55 , but it hasn't always been that way. For example, it was 2 cents when my father was a kid, and 4 cents when I was a kid. Draw the following chart. In the left hand column put 1978, 1985, 1996, 2002, 2009, and in the right hand column put $ .15 , $ .22, $ .32, $ .37, $ .44 (Don't forget to change 1978 to zero, change 1985 to 7 since it's 7 years after 1978, and change the other dates to 18, 24, 31 for the same reasons).

(a) Hit the roughly 40 buttons that I talked about in e-mail "P" and find the EXPONENTIAL regression equation to two decimal places. Also find R to six decimal places. (Nick and Zyann, due to your special situation, put that your answer is y = .17 times 1.05 to the x power. That's not what the other students will get, but I don't want to just tell them the answer, but it's what I want "you" two to put)

(b) What was the price in 1971 ? Hint- you'll need a "negative" exponent since that's "before" 1978 ( Nick and Zyann, you could do .17 divided by 1.05 divided by 1.05 etc until you use 7 of those 1.05's)

(c) What was the price in 1993 ? (Nick and Zyann, since it's a "positive" exponent, keep multiplying this time instead of dividing)

(d) What was the price in 2017, according to your formula ? (keep plugging away Nick and Zyann)

(e) In what year will the price reach $ 1.75 , obviously make two graphs and find the intersection, then add that on to 1978 (Nick and Zyann, keep mult till you reach 1.75 , keep track of how many 1.05's you use, and add that on to 1978)
Answer #1

Saluhon : Take 197820 and as so on. Year 0.23 0.82 1978 1985 1996 २००२ 2009 0.37 Now initial gunction take for forming an exp0.15 b 1.03532 Now, our function become. y = 0.15 1.03532) and the Growth factor R is - 6 = 14 R R = -1 + 1.03532 R = 0.03532@ Toice in 2011 e 9 = 2017-1978 = 39 y = 0.15 (1:03532) y = $0.58 Given, y = $1.75 , how we calculate for me 1.75 = 0.18 (1.0